# Propagation Delays of Ultra-High-Energy Cosmic Rays from Active Galactic Nuclei

### v2

@@ Abstract
We investigate the time delay incurred during ultra-high energy cosmic ray (UHECR) propagation over cosmological distances and its potential impact on the correlation between UHECR directions of arrival and sources such as Active Galactic Nuclei (AGNs), the UHECR chemical composition, and extragalactic magnetic field constraints. We propagate particles in different magnetic field configurations, spanning over an extended range of particle Larmor radii and magnetic field coherence lengths, also including attenuation losses. We find that UHECR delays could easily be comparable to (and longer than) AGN duty cycles, effectively erasing the correlation between known AGNs and UHECR anisotropies. We finally consider how strong constraints on the chemical composition of the heaviest UHECRs could enable a better characterization of extragalactic magnetic fields.
@@ 1 Introduction
Pinpointing the sources of the Ultra-High-Energy Cosmic Rays (UHECRs) may involve finding a correlation between their time/directions of arrival and the distribution of potential sources. The correlation with transient or short-lived sources, such as gamma-ray bursts and newly-born millisecond neutron stars, is already washed away by the delay induced by propagation in our Galaxy, which is likely of the order of $\sim$1000 yr (e.g., Miralda-Escudé and Waxman, 1996; Murase, 2009; Kalli, S. et al., 2011; Takami and Murase, 2012; Harari et al., 2021; van Vliet et al., 2021). For Active Galactic Nuclei (AGNs), correlation studies have mainly focused on flaring phases that would coincide with UHECR acceleration (e.g., Biermann and Strittmatter, 1987; Dermer et al., 2009; Murase et al., 2012; Farrar and Gruzinov, 2009), but little attention has been paid to analyzing propagation delays that could be compared with their much longer duty cycles.
In this work, we consider a large range of possible properties of the magnetic fields between sources and Earth to examine potential UHECR delays during propagation, and their impact on detected UHECR chemical composition. We find that UHECR propagation can introduce delays comparable to AGN duty cycles, rendering statistical correlations and anisotropy studies involving AGNs quite challenging; moreover, these conclusions likely extend to starburst galaxies, since the periods of enhanced star formation could be comparable to AGN duty cycles, for instance when fed by major galaxy mergers. Star formation can be quenched either by AGN or stellar feedback; in the former case there is a direct physical connection between the two duty cycles, in the latter quenching should kick in when massive stars go supernovae, which is still in the Myr ballpark. Within the order of magnitude uncertainty that applies to AGN duty cycles, we conclude that the arguments in this paper should apply to scenarios in which UHECR sources are either starburst galaxies or AGNs.
@@ 1.2 Magnetic Fields and UHECR Propagation
Despite the substantial body of work that has attempted to constrain the properties of $B$-fields in galaxies, galaxy clusters and beyond, their basic properties in regions that separate us from UHECR sources remain poorly known.
Previous studies have investigated the role of propagation delays in shaping the observed UHECRS flux and features. Matthews et al. (2018) examined whether past activity in nearby AGNs, Centaurus A and Fornax A, could explain the present-day UHECR flux, accounting for a possible efficient past AGN activity and the time delays accumulated during propagation. Eichmann et al. (2022); Eichmann and Kachelrieß (2023) analyzed how source lifetimes and magnetic deflections affect spectral hardening and composition evolution, highlighting the importance of extended propagation times in shaping observed spectra. In the context of the Local Supercluster, Mollerach and Roulet (2019, 2022) explored the effects of CR diffusion including finite source ages, emphasizing their influence on spectral features, time delays, and anisotropy. Adding to this picture, Fargion et al. (2018) estimated that AGN flare episodes are unlikely to correlate with UHECR arrival times, given potential propagation delays reaching up to $10^{6}$ years. In general, even relatively small delays ($\lesssim 1$Myr) make associations of transient activity and UHECR detections very challenging.
Collectively, these works have deepened our understanding of how magnetic fields, source characteristics, and propagation effects modulate UHECR observables.
Our work extends these efforts by developing a generalized framework for propagation delays that, crucially, incorporates the effects of the most updated theory of CR scattering and can be applied to several different environments: inside/around the sources, in the host galaxy and cluster, and on cosmological distances. In addition, we leverage UHECR chemical composition to place tighter constraints on the properties of extragalactic magnetic fields. Specifically, we attempt to bracket our ignorance by propagating particles —without assuming diffusion a priori— in a large range of $B$-field strengths, spectra, and coherence lengths, also considering the effects of particle losses due to photomeson and photodisintegration interactions with the cosmic microwave background (CMB), cosmic infrared background (CIB), and cosmic optical background (COB) (Dole et al., 2006). Our goals are to compare the ensuing delays with expected AGN duty cycles to estimate: i) the significance of UHECR delay and deflection angles, and ii) the effects of such delays on losses and their repercussions on the detected UHECR chemical composition.
@@ 2 Particle propagation
We consider a simple setup where test particles are propagated with a Boris pusher (e.g., Birdsall and Langdon, 1991) with different initial pitch angles $\mu$ in a prescribed magnetic field $\mathbf{B}$. A rigidity-dependent time step is chosen to ensure proper resolution of particle gyrations, with more than five steps per gyroperiod. While the magnitude of the magnetic field, $B$, remains constant during particle propagation, a turbulent component is added to alter its direction. The field direction changes by an angle $\theta$, following a random walk every coherence length $\ell_{\rm c}$, which characterizes the local magnetic turbulence. The angle $\theta$ is drawn from a Gaussian distribution centered at zero with a standard deviation of $\pi/2$, simulating sharp, strong magnetic field bends ($\delta B/B\sim 1$). The propagated particles are representative of different species, since propagation is dependent on rigidity only (except for losses).
Particle transport has been extensively studied in turbulent boxes with magnetohydrodynamical (MHD) simulations (e.g., Beresnyak et al., 2011; Cohet and Marcowith, 2016; Dundovic et al., 2020), but this approach is unfeasible for the vast range of scales involved in UHECR propagation. Instead, frameworks where particle diffusion is simulated on synthetic fields are routinely adopted to study UHECR transport (e.g., CRPropa, Alves Batista et al., 2016). Our approach, which does not assume diffusive transport but only the random walk of the magnetic field lines, allows us to explore a wide range of Larmor radii $\mathcal{R}$ much smaller/larger than the scattering mean free path (not possible in a diffusive approximation), the field coherence length $\ell_{\rm c}$, and distance to the source. Though it ignores the actual cosmological distribution of matter in the Universe, our quite simple framework is sufficient to investigate the effects of typical strengths and coherence lengths on the propagation delays, obtained by comparing the UHECR travel time $t$ with the ballistic time $t_{b}$ from the source, i.e., the photon (or neutrino) travel time.
We now consider different transport regimes, depending on whether $\mathcal{R}\gtrless l_{c}$.
@@ 4.3 Propagation Inside Galaxy Clusters
The left panel of Figure 2 shows the delays expected on a distance of 1 Mpc, the typical size of a galaxy cluster, in particular for the typical cluster values of $B\lesssim 0.1-1\rm\mu G$ and $\ell_{\rm c}\sim 10^{2}-10^{4}\rm pc$ (magenta box); such delays can already be comparable to or greater than the duty cycles of AGNs (see also Berezinsky et al., 1997; Blasi and Olinto, 1998). This is particularly important since powerful jetted AGNs are predominately found in galaxy clusters (e.g., Begelman et al., 1984; Best et al., 2007; Fang and Murase, 2018), and are prominent UHECR source candidates (e.g., Caprioli, 2015; Mbarek and Caprioli, 2019; Katz et al., 2009; Murase and Fukugita, 2019; Jiang et al., 2021). Also, over these timescales, a galaxy could change its identity and flip between AGN, starburst, or quiescent states, thus affecting any study that seeks a correlation with a specific class.

### v3

# Revisiting Propagation Delays of Ultra-High-Energy Cosmic Rays from Long-lived Sources

@@ Abstract
We revisit the time delay incurred during ultra-high energy cosmic ray (UHECR) propagation over cosmological distances and its potential impact on the correlation between UHECR directions of arrival and long-lived sources (i.e., with duty cycles of order of Myr, such as Active Galactic Nuclei (AGNs) and starburst galaxies), the UHECR chemical composition, and extragalactic magnetic field constraints. We propagate particles in different magnetic field configurations, spanning over an extended range of particle Larmor radii and magnetic field coherence lengths, also including attenuation losses. We conclude that UHECR delays could easily be comparable to (and longer than) AGN duty cycles, effectively erasing the correlation between known AGNs and UHECR anisotropies. We finally consider how strong constraints on the chemical composition of the heaviest UHECRs could enable a better characterization of extragalactic magnetic fields.
@@ 1 Introduction
Pinpointing the sources of the Ultra-High-Energy Cosmic Rays (UHECRs) involves finding a correlation between their time/directions of arrival and the distribution of potential sources. The correlation with transient or short-lived sources, such as gamma-ray bursts and newly-born millisecond neutron stars, is washed away by the delay induced by extragalactic propagation (e.g., Miralda-Escudé and Waxman, 1996; Murase and Takami, 2009; Kalli, S. et al., 2011; Takami and Murase, 2012; Harari et al., 2021; van Vliet et al., 2021), and by the delay accumulated in our Galaxy, which is likely $\gtrsim$1000 yr for particles with rigidities of $10^{19}$ or less (e.g., Alvarez-Muniz et al., 2002; Murase and Takami, 2009; Takami and Murase, 2012).
For Active Galactic Nuclei (AGNs), the same arguments apply to UHECRs produced during flaring phases (e.g., Biermann and Strittmatter, 1987; Dermer et al., 2009; Murase et al., 2012; Farrar and Gruzinov, 2009), but comparing propagation delays with their much longer duty cycles (of the order or Myr or more, as outlined below) requires modeling extra-galactic propagation.
In this work, we consider a large range of possible properties of the magnetic fields in/around UHECR sources, in galaxy clusters, and on cosmological scales to examine potential UHECR delays during propagation, and their impact on the detected UHECR chemical composition. We find that UHECR propagation generally introduces delays comparable to AGN duty cycles, rendering statistical correlations and anisotropy studies involving these sources quite challenging; moreover, these conclusions likely extend to starburst galaxies, since the periods of enhanced star formation could be comparable to AGN duty cycles, for instance when fed by major galaxy mergers. Star formation can be quenched either by AGN or stellar feedback; in the former case there is a direct physical connection between the two duty cycles, in the latter quenching should kick in when massive stars go supernovae, which is still in the Myr ballpark. Within the order of magnitude uncertainty that applies to AGN duty cycles, we conclude that the arguments in this paper should apply to scenarios in which UHECR sources are either starburst galaxies or AGNs.
@@ 1.2 Magnetic Fields and UHECR Propagation
Despite the substantial body of work that has attempted to constrain the properties of $B$-fields in galaxies, galaxy clusters and beyond, their basic properties in regions that separate us from UHECR sources remain poorly known. Previous studies have investigated the role of propagation delays in shaping the observed UHECR flux and features. Matthews et al. (2018) examined whether past activity in nearby AGNs, Centaurus A and Fornax A, could explain the present-day UHECR flux, accounting for a possible efficient past AGN activity and the time delays accumulated during propagation. Eichmann et al. (2022); Eichmann and Kachelrieß (2023) analyzed how source lifetimes and magnetic deflections affect spectral hardening and composition evolution, highlighting the importance of extended propagation times in shaping observed spectra. In the context of the Local Supercluster, Mollerach and Roulet (2019, 2022) explored the effects of CR diffusion including finite source ages, emphasizing their influence on spectral features, time delays, and anisotropy. Adding to this picture, Fargion et al. (2018) estimated that AGN flare episodes are unlikely to correlate with UHECR arrival times, given potential propagation delays reaching up to $10^{6}$ years. In general, even delays as small as $\lesssim 1$Myr make associations of transient activity and UHECR detections very challenging.
Collectively, these works have deepened our understanding of how magnetic fields, source characteristics, and propagation effects modulate UHECR observables. Numerical codes, such as SimProp (e.g., Aloisio et al., 2012) and CRPropa (e.g., Batista et al., 2015; Alves Batista et al., 2022), have been developed to study UHECR propagation. In this work we develop our own agile framework for calculating propagation delays, which incorporates the effects of the most updated theory of CR scattering and can be applied to several different environments: inside/around sources, within host galaxies and clusters, and across cosmological distances. Specifically, we attempt to bracket current uncertainties by propagating particles without assuming diffusion a priori, exploring an unprecedentedly-wide range of magnetic field strengths, spectral indices, and coherence lengths. We also account for energy losses due to photomeson production and photodisintegration from interactions with the cosmic microwave, infrared, and optical backgrounds (Dole et al., 2006).
While earlier studies have examined UHECR time delays in more restricted scenarios with specific assumptions and/or considering a limited region of the parameter space (e.g., Bhattacharjee and Sigl, 2000; Das et al., 2008; Davoudifar et al., 2011), our overarching goal is to assess whether UHECR propagation is generally subject to delays that exceed also long-lived sources, such as AGN and starburst galaxies, with their multi-Myr duty cycles. In addition, we show how a precise measurement of the UHECR chemical composition can place tighter constraints on the properties of extragalactic magnetic fields.
@@ 2 Particle propagation
We consider a simple setup where test particles are propagated with a Boris pusher (e.g., Birdsall and Langdon, 1991) with different initial pitch angles $\mu$ in a prescribed magnetic field $\mathbf{B}$. A rigidity-dependent time step is chosen to ensure proper resolution of particle gyrations, with more than five steps per gyroperiod. While the magnitude of the magnetic field, $B$, remains constant during particle propagation, a turbulent component is added to alter its direction. The field direction changes by an angle $\theta$, following a random walk every coherence length $\ell_{\rm c}$, which characterizes the local magnetic turbulence. The angle $\theta$ is drawn from a Gaussian distribution centered at zero with a standard deviation of $\pi/2$, simulating sharp, strong magnetic field bends ($\delta B/B\sim 1$). The propagated particles represent different species, since propagation is dependent on rigidity only (except for losses).
Particle transport has been extensively studied in turbulent boxes with magnetohydrodynamical (MHD) simulations (e.g., Beresnyak et al., 2011; Cohet and Marcowith, 2016; Dundovic et al., 2020), but this approach is unfeasible for the vast range of scales involved in UHECR propagation. Instead, frameworks where particle diffusion is simulated on synthetic fields are routinely adopted to study UHECR transport (e.g., CRPropa, Alves Batista et al., 2016). Our approach, which does not assume diffusive transport but only the random walk of the magnetic field lines, allows us to explore a wide range of Larmor radii $\mathcal{R}$ much smaller/larger than the scattering mean free path (not possible in a diffusive approximation), $B$ coherence length $\ell_{\rm c}$, and distance to the source. Though it ignores the actual cosmological distribution of matter in the Universe, our quite simple framework is sufficient to investigate the effects of typical strengths and coherence lengths on the propagation delays, obtained by comparing the UHECR travel time $t$ with the ballistic time $t_{b}$ from the source, i.e., the photon (or neutrino) travel time. We now consider different transport regimes, depending on whether $\mathcal{R}\gtrless l_{c}$.
@@ 4.3 Propagation Inside Galaxy Clusters
The left panel of Figure 2 shows the delays expected on a distance of 1 Mpc, the typical size of a galaxy cluster, in particular for the typical cluster values of $B\lesssim 0.1-1\rm\mu G$ and $\ell_{\rm c}\sim 10^{2}-10^{4}\rm pc$ (magenta box); such delays can already be comparable to or greater than the duty cycles of AGNs (see also Berezinsky et al., 1997; Blasi and Olinto, 1998; Matthews and Taylor, 2021; Condorelli et al., 2023). This is particularly important since powerful jetted AGNs are predominately found in galaxy clusters (e.g., Begelman et al., 1984; Best et al., 2007; Fang and Murase, 2018), and are prominent UHECR source candidates (e.g., Caprioli, 2015; Mbarek and Caprioli, 2019; Katz et al., 2009; Murase and Fukugita, 2019; Jiang et al., 2021). Also, over these timescales, a galaxy could change its identity and flip between AGN, starburst, or quiescent states, thus affecting any study that seeks a correlation with a specific class.
Propagation in galaxy clusters is most likely diffusive due to the presence of relatively strong magnetic fields, up to $\sim 1\rm\mu G$, and coherence lengths on the scales of galaxy encounters (10-100kpc). Within these environments, intermittent magnetic structures may play an important role, as they can significantly increase particle travel times, as described above (also see Condorelli et al., 2023, and references therein). As a result, heavier nuclei are more likely to undergo frequent photodisintegration, leading to substantial changes in their chemical composition. We explore the effects of photodisintegration in more detail below.